OpenStudy (anonymous):

Suzy uses a credit card to buy a $ 2000 LCD TV at 18% ,if the interest is compounded monthly , and she make no payments , how much will end up owing after 2 years ? (When does this type of thing happen)

OpenStudy (anonymous):

@ 18% compound interest Suzy, would have to pay $2000 * 18% = $360 on the first month. On the second month, That's 18 % * ($2000 * 18%) = (118%) * (118%) * $2000 = $2784.80 In General Suzy, has to pay $2000 * (118%) ^ no of months unpaid. In 24 months, $2000 * (118%) ^ 24 = $2000 * 5310.9% = $106,218.01 Thats the power of compound interest, make small things big.

OpenStudy (anonymous):

great,ty